OpenStudy (anonymous):

Solve 20x2 + 70x + 20 = 0. Answer x equals the quantity of 70 plus or minus the square root of 33 all over 4 x equals the quantity of negative 70 plus or minus the square root of 33 all over 4 x equals the quantity of negative 7 plus or minus the square root of 33 all over 4 x equals the quantity of 7 plus or minus the square root of 6 all over 4

OpenStudy (anonymous):

20x^2 + 70x + 20 = 0 10(2x^2 + 7x + 2) = 0 \[x = \frac{-b ± \sqrt{b^{2} - 4ac}}{2a}\] where a = 2, b = 7, and c = 2

OpenStudy (anonymous):

Okay, then solve?

OpenStudy (anonymous):

I'm not so good at that part

OpenStudy (anonymous):

Yes. Plug in the values and that should lead you to the answer.

OpenStudy (anonymous):

\[x = \frac{-7 ± \sqrt{7^{2} - 4(2)(2)}}{2(2)}\] Can you simplify and tell me what you get?

OpenStudy (anonymous):

Yes, one sec.

OpenStudy (anonymous):

\[-7 \pm \sqrt{33} \div 4\]

OpenStudy (anonymous):

well over 4

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Yup. Make sure that the 4 is under all of -7 ± √33

OpenStudy (anonymous):

This such equal answer # 3

OpenStudy (anonymous):

Yup.

OpenStudy (anonymous):

Yes, I just couldn't figure out how to do that . lol Thank you

OpenStudy (anonymous):

Anytime :)